Over the next three years, a firm is expected to earn economic profit of $200,000 in the first year, $300,000 in the second year, and $250,000 in the third year. After the end of the third year, the firm goes out of business. If the risk-adjusted discount rate is 9 percent for each of the next three years,

The firm can be sold today for a price of $______________.

the value of the firm is $______________.
Answer #1

Using formula for present value of series,

Present value of a firm,PV=\frac{FV1}{1+i} + FV 2 (1 + i)2 + FV3 (1+i)3

FV-Future value.

FV1=$200000 corresponding to 1st year

FV2=$300000,2nd year

FV3=$250000 ,3rd year

i=9%=.09

Therefore PV=| 200000 (14.09) + | 300000 (14.0912 + 250000 (1 +.09)3

=183486.23+252503.99+193045.87=$629036.09

Therefore value of this firm is $629036.09

The firm can be sold today for a minimum price of $629036.09